# Adaptive Tool Use in Large Language Models with Meta-Cognition Trigger

### v1

@@ 4 Benchmark-MeCa
Table 1: Performance Comparison between Naive, $P_{\text{Yes}}$ and MeCo on Metatool. Note that we are unable to calculate $P_{\text{Yes}}$ or detect the internal states of proprietary LLMs such as GPT-4-turbo.
@@
Table 2: Performance Comparison between Naive, $P_{\text{Yes}}$ and MeCo on MeCa-Tool before fine-tuning. "With context" and "without context" are abbreviated as w/ ctx and w/o ctx for brevity.
@@
Table 3: Performance Comparison between Naive, $P_{\text{Yes}}$ and MeCo on MeCa-Tool after fine-tuning on the crafted SFT data. "With context" and "without context" are abbreviated as w/ ctx and w/o ctx for brevity.
Beyond tool use, MeCa-RAG also evaluates adaptive RAG—determining whether a query can be answered directly by the LLM or requires external retrieval. RAG can be seen as a special case of tool use, where the LLM’s internal knowledge is insufficient and necessitates using a search engine to access external data. MeCa-RAG consists of:
  • Positive RAG: Cases where retrieval is essential for answering complex queries or those requiring up-to-date information.
  • Negative RAG: Cases where LLMs can answer using its internal knowledge without retrieval.
To curate the MeCa-Tool dataset, we employ a meticulous approach that began with collecting diverse scenarios from various online corpora, ensuring our synthetic APIs and conversations were grounded in realistic contexts. Based on these scenarios, we designed 500 distinct synthetic APIs emulating real-world applications across various domains and generated queries by randomly sampling from this API pool—queries that either require tool invocation, rely on the LLM’s internal knowledge, or expose cases where the provided APIs are insufficient—followed by rigorous human verification for accuracy. For the MeCa-RAG dataset, we selected common fact-based data from the RepE (Zou et al., 2023) dataset (e.g., “The Earth orbits the Sun”) to generate negative RAG queries that do not require retrieval, while positive RAG queries were generated from crawled recent news unseen by LLMs, necessitating retrieval of new information.
By incorporating both tool usage and adaptive RAG, MeCa serves as a robust benchmark for assessing LLM decision-making in complex scenarios. Detailed benchmark statistics and creation methodology for MeCa-RAG can be found in Appendix A.
5 Experiment Setup
Baselines: We evaluate MeCo against two baselines: Naive and $P_{\text{Yes}}$. The Naive baseline determines "Yes" or "No" based solely on the first token generated by the LLM, where "Yes" represents a positive indication, i.e., requiring external tools, and vice versa. The $P_{\text{Yes}}$ baseline refines this approach by computing a Yes-score,
[…]
Backbone LLMs: We employ Llama-3-8B-Instruct, Mistral-7B-Instruct-v0.3, and Llama-3-70B-Instruct as backbone models to evaluate MeCo. For brevity, we refer to them as LM3-8B, Mist-7B, and LM3-70B. Additionally, we fine-tune these models on a dataset of 4,000 tool-use query-response pairs generated by GPT-4-turbo, denoting the fine-tuned versions as LM3-8B-sft, Mist-7B-sft, and LM3-70B-sft.
Evaluation: The primary evaluation metric is decision accuracy—whether the model correctly identifies when external tools/RAG are genuinely needed. Additional metrics (precision, recall, etc.) are analyzed in Appendix C.
Prompting Strategies: We experimented with various prompting strategies, including "Yes/No" with or without explanations and Chain-of-Thought (CoT; Wei et al. (2022)). The best performance was achieved using the "Yes/No + Explanation" strategy, which is used throughout this paper. Detailed results for different prompting strategies are available in Appendix C.
Moreover, we employ two types of prompts in our experiments: 1) prompts with context, which include specific reasons why LLMs may require external tools to complete user tasks, plus five randomly sampled examples to assist the model in making decisions; and 2) prompts without context, a concise version containing only the instruction and query. Exact prompts and settings are detailed in Appendix D.
6 Experiments
We conduct extensive experiments to empirically reveal the effectiveness of MeCo on two benchmarks: Metatool and MeCa. Specifically, we evaluate MeCo in adaptive tool use on both Metatool and MeCa and in adaptive RAG on MeCa.
@@ 6.1 MeCo in Adaptive Tool Use
First, we present the distribution of meta-cognition scores collected from the pre- and post-fine-tuning models in Figure LABEL:fig:metacog_distribution. We compare the meta-cognition scores for correct and incorrect responses and visualize how these scores differentiate between correct and incorrect Yes/No answers. Our key observations and interpretations are as follows:
  • Clear Gap in Meta-Cognition Scores: In both pre- and post-fine-tuning experiments, there is a noticeable gap between the meta-cognition scores of correct and incorrect responses. Our decision-making strategy can identify and leverage this gap to distinguish between correct and incorrect Yes/No answers.
  • Higher Scores for Correct Yes, Lower for Correct No: Correct Yes responses generally have higher meta-cognition scores than incorrect Yes responses, while correct No responses have lower scores than incorrect No responses. This occurs because the meta-cognition score for Yes/No tokens depends on the token embedding. Therefore, the meta-cognition scores of different tokens are not directly comparable; the score of Yes should only be compared to other Yes scores, and the score of No should only be compared to other No scores.
@@
In our experiment, we sampled a subset of queries from the Metatool benchmark to create training data for determining the optimal thresholds for $P_{\text{Yes}}$ and MeCo. We then applied these thresholds to the test queries in both Metatool and MeCa-Tool (Task1 and Task4). Due to the fundamental difference between the queries in Metatool and those in Task 2, Task 3, Task 5, and Task 6 in MeCa-Tool, we randomly sample 100 queries from each of these categories to serve as hold-out testing data. We fit the thresholds for both $P_{\text{Yes}}$ and MeCo using the remaining data. The complete evaluation results are summarized in Table 1, Table 2, and Table 3. We make two key observations:
@@
1. Superiority of MeCo: On both benchmarks, MeCo significantly enhances the model’s naive decision accuracy regarding tool use, outperforming $P_{\text{Yes}}$ by a considerable margin, indicating the effectiveness of the meta-cognition-based trigger mechanism. Notably, MeCo’s superiority is consistent across multiple backbone models and various evaluation settings, including both with and without context, as well as pre- and post-fine-tuning.
@@
Importantly, the improvement achieved with MeCo incurs minimal costs, as it involves a fine-tuning-free and easy-to-integrate module. Note that fine-tuning and MeCo are two orthogonal approaches, and MeCo can provide additional benefits to fine-tuned models. Moreover, fine-tuned models do not transfer well to "out-of-distribution" testing scenarios. For instance, we observed performance degradation in the fine-tuned LM3-8B on Task 2 and Task 3 of MeCa-Tool. In contrast, the improvement brought by MeCo is consistent and robust across various testing scenarios.
@@
MeCo’s superiority on MeCa is particularly promising and significant. MeCa contains more complex and realistic queries and user-assistant interactions, closely mimicking real-world scenarios. This underscores MeCo’s applicability to real-world LLMs, highlighting its potential for practical deployment and effectiveness in diverse and realistic scenarios.
@@
2. Transferability: The results of Task1 and Task4 in Table 2 and Table 3 indicate that $P_{\text{Yes}}$ and MeCo, when fitted on one benchmark, can effectively transfer to other benchmarks. It’s worth noting that Metatool and MeCa feature different tool sources and styles of queries. We hypothesize that the model’s internal cognition is model-dependent, and once fitted on one benchmark, the decision strategy (i.e., the thresholds) can be transferred to other testing datasets. Although it is always better to align the decision strategy with real testing data, MeCo demonstrates satisfactory performance even when directly transferred, highlighting its robustness and adaptability.
Model Method Accuracy
@@
GPT-3.5-Turbo - 86.0
GPT-4-Turbo - 84.0
Table 4: Performance Comparison between Naive, $P_{\text{Yes}}$ and MeCo on MeCa-RAG.
@@ 6.2 MeCo in Adaptive RAG
We further evaluate the effectiveness of MeCo in the adaptive RAG task, where the LLMs need to determine whether or not to retrieve external information to address the user query. Typically, no reasons or examples are provided to the LLMs in adaptive RAG, and we follow this setting by providing no context in the prompts. The results in Table 4 further validate the effectiveness of MeCo in the adaptive RAG task, demonstrating its robustness as a trigger mechanism across various applications. Note that GPT-4-turbo has more up-to-date information and thus does not perform RAG as often as GPT-3.5-turbo and results in a lower accuracy on our benchmark.
@@ 7 Conclusion
This paper introduces the concept of adaptive tool use to advance existing tool learning paradigms, which typically rely on external tools without discrimination to address user queries. We present MeCo, a computationally efficient plug-in module that assesses LLM’s meta-cognitive states, using a probe to detect relevant signals and inform more accurate tool-use decisions. To support evaluation, we introduce a new benchmark, MeCa, specifically designed to evaluate LLMs’ awareness of tool use as well as the necessity for retrieval. We empirically validate the effectiveness of MeCo using both the Metatool and MeCa, demonstrating significant improvements in the model’s decision-making accuracy regarding the necessity for tool use and retrieval. Our findings suggest that by integrating meta-cognition into the tool usage framework, we can enhance the operational efficiency and decision-making capabilities of LLMs across diverse contexts.
@@ 8 Limitations
We have omitted a fine-grained evaluation in the MeCa benchmark that measures the model’s end-to-end performance in correctly using tools with appropriate parameters. This includes determining whether tool use is necessary, parameter filling, and answer generation, etc. These tasks require capabilities beyond tool-use decision-making and would involve significant human effort to evaluate the final outputs. Our work mainly focuses on introducing adaptive tool use and leveraging meta-cognition for self-assessment in deciding when tool use is necessary. Although applying MeCo to parameter filling for tool use in LLMs is feasible, it also introduces additional latency. How to better apply MeCo to improve the parameter filling accuracy and the end-to-end evaluation remains an important direction for future work.
@@ A.1 MeCa Statistics
Table 5 summarizes the statistics of the MeCa. In Task1 and Task4, the positive queries require a specific external tool to address the user queries, while the negative queries require no external tools and can be solved by the LLM’s internal capabilities. In Task2 and Task5, we provide a tool name and its description along with the user query, asking the LLMs to determine whether they need to use this specific tool to address the user queries. The neutral queries in Task2 and Task3 indicate that these queries require external tools, but the provided tool is irrelevant to addressing the user query. In Task3 and Task6, we provide a list of tools (ranging from 2 to 5) along with the user query. For multi-turn queries, there is a dialogue between the user and the LLM assistant, where the assistant needs to determine whether it should rely on external tools to address the user query in the final round of the conversation.
@@
Table 5: Tool Usage Categories and Counts
We directly transfer the $l_{yes}$ and $l_{no}$ thresholds of MeCo, fitted on the Metatool dataset, to Task1 and Task4 in MeCa-Tool, and present the results in Table 2 and Table 3. Because the rest of the tasks in MeCa-Tool are very different and more complex than the user queries in MetaTool, we randomly sample 100 queries from each category in Task2, Task3, Task5, and Task6, and use these queries as the hold-out testing data. We use the remaining data to fit the thresholds for $P_{yes}$ and MeCo.
@@ A.2 MeCa Creation
To curate the MeCa-Tool dataset, we employed a meticulous and structured approach that ensures the queries are relevant to current LLM capabilities. The process unfolded as follows:
@@
    Collection of diverse scenarios: We began by gathering a broad spectrum of domains and conversational scenarios from various online corpus. This initial step ensures that the subsequent generated synthetic APIs and conversations are grounded in realistic and diverse settings.
@@
    Synthetic APIs design: Leveraging the collected scenarios, we then synthetically design 500 distinct APIs by emulating examples found in real-world applications, ensuring that they span multiple domains.
@@
    Query generation: For each query, APIs are randomly sampled from our synthetic APIs pool. User queries are then constructed based on sampled APIs, which may: (i) Require the invocation of the provided APIs; (ii) Not require any tool invocation, relying solely on the LLM’s internal knowledge; or (iii) Involve cases where the provided APIs does not include the necessary tools to answer the query directly.
@@
    Human Verification: After the queries were constructed, they underwent a rigorous human review process. This critical step verified the validity and correctness of the data, ensuring that each query aligns with its intended category and meets quality standards.
@@
The dataset was constructed as follows: we selected a subset of fact-based data from the RepE dataset (Zou et al., 2023), which consists of common, well-known facts, such as “The Earth orbits the Sun.” These facts were used as model responses, and the leading proprietary LLM (i.e., GPT-4-turbo) was instructed to generate corresponding user queries. Since these queries involve common knowledge that is embedded in LLMs, they do not require retrieval and thus serve as negative RAG examples. For positive RAG examples, we scraped recent news articles from the past few months, ensuring that this content has not been seen by LLMs. We then followed a similar process as mentioned above to generate user queries based on the latest information. This process resulted in queries that require retrieval as they involve knowledge that is unknown or not yet integrated into the LLM’s training data. The detailed distribution of MeCa can be found in Figure 4.
@@ Appendix B Related Work
LLMs have progressed from understanding and generating human-like text to utilizing external tools based on natural language instructions. This evolution expands their application beyond basic conversational tasks to enable dynamic interactions across diverse functional domains, such as facility management and professional services (Patil et al., 2023; Liu et al., 2023b; Qin et al., 2023; Chen et al., 2023). For example, Toolformer (Schick et al., 2024) enables LLMs to use external tools via simple APIs through a supervised fine-tuning (SFT) model. Liu et al. (2024c) demonstrate strong executable functional API calls across different domains. ToolACE (Liu et al., 2024b) trained on synthesized data, achieves state-of-the-art results on the Berkeley Function-Calling Leaderboard (Yan et al., 2024), even with a relatively small model size of 8B parameters. Despite their growing popularity and capabilities, tool use in LLMs often depends on strategies like verbal feedback, which are hampered by the quality of the datasets used for fine-tuning. Several benchmarks/datasets have been developed to support tool use in a data-centric way, such as API-Bank (Li et al., 2023), which provides a set of tool-use dialogues with various APIs to assess the LLM’s tool use capabilities, Toolalpaca (Tang et al., 2023) constructs a comprehensive tool-use corpus derived from collected real-world APIs, designed specifically to fine-tune LLMs for better tool utilization. ToolBench (Qin et al., 2023) focuses on creating a synthetic instruction-tuning dataset for tool use. However, these methods rely solely on superficial textual information, without probing deeper into the LLM’s internal states to explain or justify when and why a tool should be called, resulting in an inability to accurately determine the optimal timing for tool invocation.
@@
RAG has shown success in supporting AI systems that require up-to-date information or access domain-specific knowledge, particularly where the scope of queries is not seen in the training data of LLMs (Lewis et al., 2020; Ren et al., 2023; Vu et al., 2023; Izacard et al., 2023). This paper is also consistent with the trend of towards adaptive RAG paradigm, which is designed to assess whether a query can be directly answered by the LLMs or requires external data retrieval (Asai et al., 2023; Jiang et al., 2023). Specifically, a simple query within the LLM’s knowledge should be directly answered by the LLMs themselves. On the other hand, for complex queries or questions about data they have not been trained on, RAG intervenes to prevent incorrect out-of-date answers or hallucination (Ji et al., 2023). This mechanism allows RAG to dynamically adjust operational strategies of retrieval-augmented LLMs by assessing the boundary of LLM’s self-knowledge and the complexity of the query, thereby minimizing unnecessary computational overhead when the queries are answerable by LLMs themselves. Similar to the LLMs function-calling, the decision of retrieval timing typically hinges on three primary methods: (i) explicit verbal feedback from LLMs (Ding et al., 2024), (ii) enhancements through fine-tuning (Asai et al., 2023), or (iii) probability-based metrics (Kadavath et al., 2022; Jiang et al., 2023). Specifically, He et al. (2021) proposed enhancing the retrieval time efficiency by computing the probability of the next token via interpolating an LLM with a distribution calculated from the $k$ nearest context-token pairs. Drozdov et al. (2022) further extend $k$NN-LM to the adaptive paradigm by assigning the interpolation coefficient according to the retrieval quality measured by semantic similarity. Asai et al. (2023) introduce Self-RAG to improve generation quality and factuality by enabling adaptive retrieval and self-reflection. In contrast, this paper conceptualizes RAG as an external tool and highlights the importance of understanding the internal states of an LLM when developing the retrieval policy.
@@
However, there is a considerable discrepancy between LLM’s decision mechanisms (often based on verbalized responses) and their internal cognition (Zou et al., 2023). The internal workings of LLMs are usually unclear, and this lack of transparency poses unwanted risks in downstream decision-making. Therefore, understanding and interpreting LLMs is crucial for elucidating their behaviors and limitations. To address this challenge, various explanations that provide insights into the inner workings of LLMs have been proposed (Zhao et al., 2024): (i) Probing-based explanations: Probing uses vector representations to measure embedded knowledge (Peters et al., 2018; Jawahar et al., 2019) or examines specific knowledge during the LLM’s generation process (Li et al., 2022), (ii) Neuron-level explanation: neuron analysis identifies critical neurons that are essential for model’s performance (Antverg and Belinkov, 2021; Bills et al., 2023), (iii) representation engineering (RepE): RepE leverages techniques inspired by cognitive neuroscience to identify and enhance the transparency of LLMs by uncovering their internal cognitive states (Zou et al., 2023). In this paper, we aim to detect the internal cognition of LLMs, and intervene LLM’s decisions, i.e., ensuring more precise decisions on tool use and retrieval timing.

### v2

@@ 6.1 Reduced Tool Invocations and Latency
Table 3: Decision accuracy (%) comparison between Naive, $P_{\text{Yes}}$ and MeCo on the MetaTool benchmark. Note that we are unable to calculate $P_{\text{Yes}}$ or detect the internal states of proprietary LLMs such as GPT-4-turbo. As a result, GPT-4-turbo uses the Naive decision-making strategy.
@@
Table 4: Decision accuracy (%) comparison between Naive, $P_{\text{Yes}}$, and MeCo on MeCa-Tool before fine-tuning. "With context" and "without context" are abbreviated as w/ ctx and w/o ctx, respectively.
@@
Table 5: Decision accuracy (%) comparison between Naive, $P_{\text{Yes}}$ and MeCo on MeCa-Tool after fine-tuning on the crafted SFT data. "With context" and "without context" are abbreviated as w/ ctx and w/o ctx, respectively.
@@ 6.2 MeCo in Adaptive Tool Use
In this experiment, we sampled a subset of queries from the MetaTool benchmark to construct validation data for determining the optimal thresholds for $P_{\text{Yes}}$ and MeCo. These thresholds were then applied to the test queries in both MetaTool and MeCa-Tool (Task 1 and 4). Given the fundamental differences between the queries in MetaTool and those in Task 2, 3, 5, and 6 of MeCa-Tool, we randomly sampled 100 queries from each of these categories to serve as held-out test data. The thresholds for both $P_{\text{Yes}}$ and MeCo were fit using the remaining data in each respective category. Comprehensive evaluation results are reported in Table 3, Table 4, and Table 5. We report the decision accuracy of GPT-4-turbo as a reference point to illustrate the difficulty level of the queries in the MetaTool benchmark. We highlight two key observations:
@@
1. Superiority of MeCo: Across both benchmarks, MeCo significantly improves the model’s decision accuracy regarding tool use, outperforming the $P_{\text{Yes}}$ baseline by a notable margin. This highlights the effectiveness of the meta-cognition-based trigger mechanism. Crucially, MeCo’s advantages are consistent across different backbone models and evaluation settings—both with and without context, as well as before and after fine-tuning.
@@
Importantly, the gains from MeCo come at minimal cost: it is a fine-tuning-free, easily integrable module. Note that fine-tuning and MeCo are two orthogonal approaches, and MeCo can provide additional benefits to fine-tuned models. Notably, fine-tuned models often struggle to generalize to "out-of-distribution" scenarios. For instance, we observed a performance drop in the fine-tuned LM3-8B on Tasks 2 and 3 of MeCa-Tool. In contrast, MeCo’s improvements are robust and consistently maintained across varied testing conditions.
@@
The strong performance of MeCo on MeCa is particularly noteworthy, as MeCa presents more complex and realistic queries and user–assistant interactions that better reflect real-world usage. This underscores MeCo ’s potential for practical deployment and its effectiveness in diverse and realistic settings.
@@
2. Transferability: Results from Tasks 1 and 4 in Table 4 and Table 5 show that both $P_{\text{Yes}}$ and MeCo, when fitted on one benchmark, can generalize effectively to others. Despite the differences in tool sources and query styles between MetaTool and MeCa benchmarks, both methods demonstrate promising cross-benchmark performance. We hypothesize that an LLM’s internal meta-cognition is primarily model-dependent, and once a decision strategy (i.e., thresholds) is learned on one dataset, it can generalize to others. Although aligning the decision strategy with the target data is always preferable, MeCo performs robustly even in direct transfer settings, highlighting its robustness and adaptability.
Model Method Accuracy (%)
@@
GPT-4-Turbo Naive 84.0
Table 6: Decision accuracy (%) comparison between Naive, $P_{\text{Yes}}$ and MeCo on MeCa-RAG.
@@ 6.3 MeCo in Adaptive RAG
We further evaluate the effectiveness of MeCo in the adaptive RAG task, where LLMs must decide whether to retrieve external information to accurately answer a user query. In line with standard settings for adaptive RAG, no explicit reasons or examples are provided in the prompts—only the raw query is given. To contextualize the difficulty of the queries in MeCa-RAG, we report the decision accuracy of GPT-4-turbo as a reference. As shown in Table 6, MeCo consistently demonstrates robust performance across multiple models and outperforms the baseline methods, validating its utility as a general-purpose trigger mechanism beyond tool use.
@@ 7 Conclusion
We introduce the concept of adaptive tool use, which advances traditional tool-learning paradigms that often invoke external tools indiscriminately. We propose MeCo, a lightweight and fine-tuning-free plug-in module that leverages meta-cognitive signals within LLMs to make more informed decisions about tool invocation. MeCo uses a probing mechanism to assess internal representations, enabling more precise determinations of when external tools and retrieval are truly necessary. To support rigorous evaluation, we introduce MeCa, a new benchmark that captures both tool-use and retrieval-awareness across realistic and diverse scenarios. Through extensive experiments on both the MetaTool and MeCa benchmarks, we demonstrate that MeCo significantly enhances decision accuracy regarding the necessity of external tools and retrieval, across multiple models and settings. Our findings suggest that incorporating meta-cognition into the LLM tool-use framework can yield more efficient and reliable decision-making in practical applications.
@@ 8 Limitations
Our current evaluation of the MeCa benchmark primarily focuses on the decision-making component—whether tool use or retrieval is necessary—rather than the full end-to-end performance of LLMs in completing tasks using tools. This includes downstream tasks such as selecting the correct tool, filling in parameters appropriately, and generating accurate final responses. Evaluating these stages would require substantial human annotation effort and introduces complexity beyond the current scope. While MeCo could be extended to assist in downstream tasks such as parameter filling, doing so would increase inference latency. Exploring how to effectively integrate MeCo into the full tool-use pipeline—balancing accuracy, efficiency, and usability—remains an important avenue for future research.
@@ A.1 MeCa Statistics
Table 7 summarizes the statistics of MeCa. In Tasks 1 and 4, positive queries require a specific external tool to address the user queries, while negative queries can be answered using the LLM’s internal capabilities alone. In Tasks 2 and 5, a tool name and its description are provided alongside the user query; the LLM must determine whether this specific tool is necessary to resolve the query. Neutral queries in Tasks 2 and 3 indicate that external tools are needed, but the given tool is irrelevant to solving the query. In Tasks 3 and 6, a list of tools (ranging from two to five) is provided with the user query. For multi-turn queries, a dialogue between the user and the assistant is presented, and the assistant must decide whether to rely on external tools to respond to the final turn.
@@
Table 7: Tool use categories and counts.
We directly transfer the $l_{yes}$ and $l_{no}$ thresholds of MeCo, fitted on the MetaTool dataset, to Tasks 1 and 4 in MeCa-Tool, with results shown in Table 4 and Table 5. Since the remaining tasks in MeCa-Tool (Tasks 2, 3, 5, and 6) differ significantly from MetaTool and involve more complex user queries, we randomly sample 100 queries from each of these tasks as hold-out test data. The remaining data is used to fit the thresholds for $P_{\text{yes}}$ and MeCo.
@@ A.2 MeCa Creation
To construct the MeCa-Tool dataset, we followed a meticulous and structured process to ensure the queries are aligned with current LLM capabilities:
@@
    Collection of diverse scenarios: We began by gathering a wide range of domains and conversational scenarios from various online corpora. This step ensured that the subsequent synthetic APIs and conversations were grounded in realistic and diverse contexts.
@@
    Synthetic API design: Based on the collected scenarios, we synthetically created 500 distinct APIs, emulating patterns from real-world applications to span multiple domains.
@@
    Query generation: For each query, APIs were randomly sampled from the synthetic API pool. User queries were then constructed to reflect one of the following types: (i) Queries that require invoking the provided APIs; (ii) Queries solvable without external tools, relying solely on the LLM’s internal knowledge; or (iii) Queries where the provided APIs do not include the necessary tools to directly answer the query.
@@
    Human verification: All generated queries underwent a rigorous human review process to ensure validity, correctness, and proper categorization. This step guaranteed high-quality annotations and alignment with task definitions.
@@
The dataset was constructed as follows: we selected a subset of fact-based entries from the RepE dataset (Zou et al., 2023), which contains well-known facts such as "The Earth orbits the Sun." These facts were used as target responses, and the leading proprietary LLM (i.e., GPT-4-turbo) was prompted to generate corresponding user queries. Since these queries involve widely known information already embedded in LLMs, they do not require retrieval and thus serve as negative RAG examples. For positive RAG examples, we scraped recent news articles from the past few months, ensuring that the content was unlikely to have been seen during LLM training. Using a similar prompting process, we generated user queries based on this up-to-date information. These queries require retrieval, as they concern facts that are either unknown to or not yet encoded in the model’s training data. The overall distribution of MeCa is illustrated in Figure 9.
@@ Appendix B Related Work
LLMs have progressed from understanding and generating human-like text to utilizing external tools based on natural language instructions. This evolution expands their application beyond basic conversational tasks to enable dynamic interactions across diverse functional domains, such as facility management and professional services (Patil et al., 2023; Liu et al., 2023b; Qin et al., 2023; Chen et al., 2023; Zhang and Lee, 2025). For example, Toolformer (Schick et al., 2024) enables LLMs to use external tools via simple APIs through a supervised fine-tuning (SFT) model. Liu et al. (2024c) demonstrates strong executable functional API calls across different domains. ToolACE (Liu et al., 2024b) trained on synthesized data, achieves state-of-the-art results on the Berkeley Function-Calling Leaderboard (Yan et al., 2024), even with a relatively small model size of 8B parameters. Despite their growing popularity and capabilities, tool use in LLMs often depends on strategies like verbal feedback, which are hampered by the quality of the datasets used for fine-tuning. Several benchmarks/datasets have been developed to support tool use in a data-centric way, such as API-Bank (Li et al., 2023), which provides a set of tool-use dialogues with various APIs to assess the LLM’s tool use capabilities, Toolalpaca (Tang et al., 2023) constructs a comprehensive tool-use corpus derived from collected real-world APIs, designed specifically to fine-tune LLMs for better tool utilization. ToolBench (Qin et al., 2023) focuses on creating a synthetic instruction-tuning dataset for tool use. However, these methods rely solely on superficial textual information, without probing deeper into the LLM’s internal states to explain or justify when and why a tool should be called, resulting in an inability to accurately determine the optimal timing for tool invocation.
@@
RAG has shown success in supporting AI systems that require up-to-date information or access domain-specific knowledge, particularly where the scope of queries is not seen in the training data of LLMs (Lewis et al., 2020; Ren et al., 2023; Vu et al., 2023; Izacard et al., 2023; Dong et al., 2025). This paper is also consistent with the trend towards the adaptive RAG paradigm, which is designed to assess whether a query can be directly answered by the LLMs or requires external data retrieval (Asai et al., 2023; Jiang et al., 2023). Specifically, a simple query within the LLM’s knowledge should be directly answered by the LLMs themselves. On the other hand, for complex queries or questions about data they have not been trained on, RAG intervenes to prevent incorrect out-of-date answers or hallucination (Ji et al., 2023). This mechanism allows RAG to dynamically adjust operational strategies of retrieval-augmented LLMs by assessing the boundary of LLM’s self-knowledge and the complexity of the query, thereby minimizing unnecessary computational overhead when the queries are answerable by LLMs themselves. Similar to the LLMs function-calling, the decision of retrieval timing typically hinges on three primary methods: (i) explicit verbal feedback from LLMs (Ding et al., 2024), (ii) enhancements through fine-tuning (Asai et al., 2023), or (iii) probability-based metrics (Kadavath et al., 2022; Jiang et al., 2023). Specifically, He et al. (2021) proposed enhancing the retrieval time efficiency by computing the probability of the next token via interpolating an LLM with a distribution calculated from the $k$ nearest context-token pairs. Drozdov et al. (2022) further extends kkNN-LM to the adaptive paradigm by assigning the interpolation coefficient according to the retrieval quality measured by semantic similarity. Asai et al. (2023) introduces Self-RAG to improve generation quality and factuality by enabling adaptive retrieval and self-reflection. In contrast, this paper conceptualizes RAG as an external tool and highlights the importance of understanding the internal states of an LLM when developing the retrieval policy.
@@
However, there is a considerable discrepancy between the LLM’s decision mechanisms (often based on verbalized responses) and their internal cognition (Zou et al., 2023). The internal workings of LLMs are usually unclear, and this lack of transparency poses unwanted risks in downstream decision-making. Therefore, understanding and interpreting LLMs is crucial for elucidating their behaviors and limitations. To address this challenge, various explanations that provide insights into the inner workings of LLMs have been proposed (Zhao et al., 2024): (i) Probing-based explanations: Probing uses vector representations to measure embedded knowledge (Peters et al., 2018; Jawahar et al., 2019) or examines specific knowledge during the LLM’s generation process (Li et al., 2022), (ii) Neuron-level explanation: neuron analysis identifies critical neurons that are essential for model’s performance (Antverg and Belinkov, 2021; Bills et al., 2023), (iii) representation engineering (RepE): RepE leverages techniques inspired by cognitive neuroscience to identify and enhance the transparency of LLMs by uncovering their internal cognitive states (Zou et al., 2023). In this paper, we aim to detect the internal cognition of LLMs, and intervene in LLM’s decisions, i.e., ensuring more precise decisions on tool use and retrieval timing.